# BoViLA: Bootstrapping Video-Language Alignment via LLM-Based Self-Questioning and Answering

### v1

@@ Abstract
The development of multi-modal models has been rapidly advancing, with some demonstrating remarkable capabilities. However, annotating video-text pairs remains expensive and insufficient. Take video question answering (VideoQA) tasks as an example, human annotated questions and answers often cover only part of the video, and similar semantics can also be expressed through different text forms, leading to underutilization of video. To address this, we propose BoViLA, a self-training framework that augments question samples during training through LLM-based self-questioning and answering, which help model exploit video information and the internal knowledge of LLMs more thoroughly to improve modality alignment. To filter bad self-generated questions, we introduce Evidential Deep Learning (EDL) to estimate uncertainty and assess the quality of self-generated questions by evaluating the modality alignment within the context. To the best of our knowledge, this work is the first to explore LLM-based self-training frameworks for modality alignment. We evaluate BoViLA on five strong VideoQA benchmarks, where it outperforms several state-of-the-art methods and demonstrate its effectiveness and generality. Additionally, we provide extensive analyses of the self-training framework and the EDL-based uncertainty filtering mechanism. The code will be made available at https://github.com/dunknsabsw/BoViLA.
Introduction
@@
Recent advances in multimodal large models (MLLMs) have demonstrated the effectiveness of scaling laws in visual instruction fine-tuning. However, the continued advancement of MLLMs is hindered by the high cost of human annotation required for visual instruction data. In fact, this supervised fine-tuning paradigm does not fully exploit the rich information available in the visual modality data or the internal knowledge of frozen large language models (LLMs), yet merely increasing the training data without optimizing data utilization can be inefficient. For example, in conventional video question-answering (VideoQA) tasks, models typically predict answers based on a given video and its associated annotated question. This approach, however, is suboptimal for effective learning. On one hand, as noted in (McQuivey 2008) that “A Video Is Worth 1.8 Million Words”, videos often contain extensive information that can be described in various forms of language. However, typical datasets offer text that is both limited in length and uniform in structure, which significantly underutilizes the rich information embedded in videos. This restricts the model’s learning to the specific annotated question-answer pairs, limiting its ability to generalize to semantically similar questions presented in different formats. Consequently, this naive training paradigm hinders the model’s capacity for analogical reasoning. On the other hand, this mechanical and passive supervised training method is notably inferior compared to human learning processes, which tends to be more active. Humans often draw upon past experiences and cognition to enrich their understanding of current events, hence proactively pose new questions and seeking answers for a more comprehensive and profound grasp of the situation. This form of learning, which integrates historical knowledge and is often abstracted as “world model”.
To address this, we introduce Bootstrapping Video-Language Alignment (BoViLA) training framework via LLM-based self-questioning and answering. It further exploits internal knowledge of LLMs and the rich information in videos. BoViLA includes two roles, the questioner and the answerer, both played by the same model and improve each other alternately through self-questioning and answering, as shown in Fig 1. Questioner generates new questions for enabling itself to further extract aligned knowledge from videos and unleash power of the LLM. Answerer provides feedback to questioner in terms of the self-generated question. This framework features efficiently employment of video data and the internal historical knowledge of LLMs.
Additionally, we also apply EDL-estimated uncertainty to filter out low-quality questions resulting from modal unalignment. We enhance the vanilla EDL by decoupling the direction and magnitude of evidence, as directly applying a non-negative activation function to the logits of an LLM, where most parameters are frozen and the logits have high dimensionality, can result in substantial information loss.
We verify the effectiveness of the BoViLA on five challenging VideoQA benchmarks: STAR(Wu et al. 2024a), How2QA(Li et al. 2020), DramaQA(Choi et al. 2021), TVQA(Lei et al. 2018), and VLEP(Lei et al. 2020), where BoViLA outperforms several strong baselines. Moreover, we present extensive ablation studies as shown in Table 3 and Appendix. To sum up, our contributions are as follows:
@@
Related Work
LLMs for multi-modal understanding
As LLMs have demonstrated impressive capabilities(Brown et al. 2020; Ouyang et al. 2022; Raffel et al. 2020; Touvron et al. 2023; Chiang et al. 2023), there has been increasing interest in exploring multi-modal language models (MLLMs)(Hu et al. 2024) with visual capabilities. Unlike the more costly joint visual-language pre-training methods, some approaches focus on training lightweight visual-language connectors that endow LLMs with visual abilities(Ma et al. 2023). These methods efficiently utilize the linguistic knowledge of LLMs and the visual knowledge of pre-trained visual encoders for cross-modal alignment.
For instance, Flamingo(Alayrac et al. 2022) utilize a cross-attention mechanism to inject visual knowledge into the LLM, which is the so-called Perceiver Resampler. LLaMA-Adapter(Zhang et al. 2023a) applies a linear projection along with prompt adaptation to incorporate visual information, effectively projecting visual embeddings into the input space of the LLM. Additionally, BLIP-2(Li et al. 2023) trains a module called the Q-former to bridge the modal gap between pre-trained visual encoders and LLMs, enhancing the model’s multi-modal understanding.
Video Question-Answering
@@
Traditional VideoQA methods relied on training separate visual and text encoders, along with temporal modeling and answering modules(Qian et al. 2023; Xiao et al. 2022; Lei et al. 2021). However, with the advent of large language models (LLMs), there is a growing trend towards using LLM-based approaches due to their advanced reasoning abilities(Yu et al. 2024; Wang et al. 2023b; Ko et al. 2023; Zhang, Li, and Bing 2023; Yu, Yoon, and Bansal 2024).
@@
For instance, SeViLA(Yu et al. 2024) uses an LLM to select keyframes from a video and employs another LLM to answer questions based on these keyframes, fully leveraging LLMs’ capabilities. VLAP(Wang et al. 2023b) improves on this by introducing a Frame-Prompter and QFormer-Distiller for more efficient modality alignment. CREMA(Yu, Yoon, and Bansal 2024) trains a multimodal Q-former to integrate information from different modalities to answer questions. We notice the recent work of LLaMA-VQA(Ko et al. 2023) as closest to ours, which trains only a linear layer and adaptor to enable LLMs to understand videos, and enhancing modality alignment through multi-task learning by reconstructing both questions and video. In contrast, our work prompts LLMs to simultaneously engage in self-questioning and answering. The major differences are that (i) We use self-generated questions as augmented training data and ask the model to answer them correctly. (ii) We further enhance the model’s questioning ability by encouraging it to answer correctly from self-generated questions. In other words, the loss of the answerer on the self-generated questions propagates gradients back to the parameters of the questioner.
LLM-Based Bootstrapping Training
Early research has extensively explored the use of LLMs to generate training data for other models, which leverages the capabilities of LLMs to reduce the dependency on human labor for data collection(Lee et al. 2024; Liu et al. 2022; Meng et al. 2023; Wang et al. 2024; Mekala et al. 2022). Recently, there has been increasing interest in using LLMs to generate data for their own training(Ulmer et al. 2024; Zhao et al. 2024; Wang et al. 2022a; Huang et al. 2022; Amini et al. 2022).
@@
For example, Wang et al. (2022a) enables models to generate data for instruction fine-tuning, focusing on data efficiency and general-purpose tasks. Huang et al. (2022) use Chain-of-Thought prompting and self-consistency to generate rationale-augmented answers without labeled data. Ulmer et al. (2024) focuses on a single improve step and employs a conceptually simpler supervised finetuning strategy instead of RL. Zhao et al. (2024) investigates how self-generation can further enhance an instruction-finetuned model’s ability to execute task-specific instructions. Our work differs from prior efforts in several key ways: (i) We focus on multi-modal tasks rather than text-only ones; (ii) We update model parameters end-to-end rather than alternately performing these processes offline to improve modality alignment; (iii) This dual approach enhances learning efficiency, offering a faster and more convenient solution.
@@
Furthermore, training data generation can be seen as a form of knowledge distillation(Lei and Tao 2023; Wang et al. 2022a; Yang et al. 2024), while our self-questioning and answering method can also be seen as self-distillation.
Uncertainty Estimation Model
@@
Numerous studies have explored models capable of estimating uncertainty to enhance reliability and trustworthiness(Zhang et al. 2021; Xiao et al. 2021; Li 2022; Izmailov et al. 2021; Gal and Ghahramani 2016; Amini et al. 2020; Sensoy, Kaplan, and Kandemir 2018). EDL(Sensoy, Kaplan, and Kandemir 2018), as one of these approaches, models ”second-order probabilities” over logits based on Dempster-Shafer Theory(Shafer 1992) and Subjective Logic(Jsang 2018) to capture uncertainty conveniently and accurately in various fields(Han et al. 2022; Huang et al. 2024a; Ma et al. 2024; Huang et al. 2024b, 2023; liu2024adaptive,chen2024toward), particularly for out-of-distribution (OOD) samples.
In this work, we leverage EDL-estimated uncertainty to evaluate the modality alignment within the context and employ it for ”soft filtering” of the model’s self-generated questions. To the best of our knowledge, we are the first to explore the integration of EDL with LLMs.
Methdology
@@
As shown in Figure 2, our model architecture consists of an LLM decoder, a learnable linear layer for mapping visual tokens to the text embedding space, a lightweight adaptor for task-specific fine-tuning, and an EDL head to estimate uncertainty. For videos, we first extract frames $\bm{v}=\{v_{1},v_{2},\cdots,v_{N_{v}}\}$ and use a pretrained visual encoder $E(\cdot)$ to extract their features $E(\bm{v})=\{E(v_{1}),E(v_{2}),\cdots,E(v_{N_{v}})\}\in{\mathbb{R}}^{{N_{v}}\times D}$. These features are then mapped to the text embedding space with the learnable linear layer $f_{\theta}(\cdot)$, and an extra learnable temporal embedding $\bm{t}=\{t_{1},t_{2},\cdots,t_{N_{v}}\}\in{\mathbb{R}}^{{N_{v}}\times D}$ is added, i.e.
@@
For text inputs such as task instructions, questions, or answers, we use the LLM’s tokenizer and token embedding module to obtain the corresponding tokens and embedding. Specifically, for questions, we get the tokens $\bm{q}=\{q_{1},q_{2},\cdots,q_{N_{q}}\}$ and their embedding $\bm{h}^{0}_{q}=\{q^{0}_{1},q^{0}_{2},\cdots,q^{0}_{N_{q}}\}$. Similarly, for answers, we obtain the tokens $\bm{a}=\{a_{1},a_{2},\cdots,a_{N_{a}}\}$ and their embedding $\bm{h}^{0}_{a}=\{a^{0}_{1},a^{0}_{2},\cdots,a^{0}_{N_{a}}\}$. Then, we concatenate the video and text tokens together as input to the LLM. As task-specific fine-tuning is necessary, we employed several prevalent PEFT methods such as LoRA(Hu et al. 2021), adapter methods(Hu et al. 2023; Zhang et al. 2023b), and prefix-tuning(Li and Liang 2021; Zhang et al. 2023b) for efficient modal alignment.
Bootstrapping Training Framework
@@
Questioner.
When the model acts as the questioner, we instruct it to generate a question based on the video, answer, and seed question, as shown in Figure 2. Assuming the LLM has $l$ layers, the hidden states from the final layer are $h^{l}_{q}=\{q^{l}_{1},q^{l}_{2},\cdots,q^{l}_{N_{q}}\}$. For gradient backpropagation, we use Gumbel-Softmax(Jang, Gu, and Poole 2016) for sampling, formulated as below:
@@
$\mathcal{L}_{\mathrm{v\overline{q}a}}=-\log p(\bm{a}|\bm{v},\bm{\overline{q}})=-\sum\nolimits_{i=1}^{N_{a}}\log p(a_{i}|\bm{v},\bm{\overline{q}},a_{<i}).\vspace{-2mm}$ (6)
Regularization Based on Seed Questions.
A key point of BoViLA is the differentiability of self-generated questions. If answerer fails to predict the target answer from $\overline{q}$, then penalty will be applied to the questioner to improve itself. In contrast, if answerer succeed in answering $\overline{q}$, that means $\overline{q}$ is considered a ”good” question by answerer. However, this can easily lead to ’information leakage’, where the questioner creates meaningless questions but contain target answers implicitly. To address this, we apply seed-based question regularization to constrain the generation space of questioner as follows:
@@
EDL-based Filter

### v2

# Uncertainty-Guided Self-Questioning and Answering for Video-Language Alignment

@@ Abstract
The development of multi-modal models has been rapidly advancing, with some demonstrating remarkable capabilities. However, annotating video-text pairs remains expensive and insufficient. Take video question answering (VideoQA) tasks as an example, human annotated questions and answers often cover only part of the video, since the corresponding text is often short and monotonous, leading to underutilization of video. To address this, we propose a Bootstrapping Video-Language Alignment framework (BoViLA), a self-training method that augments question samples during training process through LLM-based self-questioning and answering, which help model exploit video information and the internal knowledge of LLMs more thoroughly to improve modality alignment. However, low-quality self-generated questions may instead contaminate the performance, especially in the early stages of training, as we have observed in our experiments. To filter bad self-generated questions, we introduce Evidential Deep Learning (EDL) to estimate uncertainty and assess the quality of self-generated questions by evaluating the modality alignment within the context. To the best of our knowledge, this work is the first to explore LLM-based self-training frameworks for modality alignment. We evaluate BoViLA on five strong VideoQA benchmarks, where it outperforms several state-of-the-art methods and demonstrate its effectiveness and generality. Additionally, we provide extensive analyses of the self-training framework and the EDL-based uncertainty filtering mechanism. The code will be made available.
keywords:
Self-questioning and answering , Large language models , Evidential deep learning , Modality alignment
[…]
[a] organization=TeleAI, city=Beijing, postcode=100007, country=China
[…]
[b] organization=School of Mathematics and Statistics, addressline=Xi’an Jiaotong University, city=Xi’an, postcode=710049, country=China
[…]
[c] organization=The University of Hong Kong, addressline=Pokfulam Road, city=Hong Kong, postcode=999077, country=China
{highlights}
Low data utilization and inefficient modality alignment in existing video-text methods motivate a novel self-training framework.
A self-questioning and answering training environment is constructed to enhance video-language alignment.
Evidential Deep Learning (EDL) is improved and successfully applied to LLMs for uncertainty estimation.
An EDL-based uncertainty-guided mechanism is devised to assess and filter self-generated questions.
The framework achieves higher data efficiency and improved modality alignment without additional human annotations.
1 Introduction
Recent advances in multimodal large models (MLLMs) have demonstrated the effectiveness of scaling laws in visual instruction fine-tuning. However, the continued advancement of MLLMs is hindered by the high cost of human annotation required for visual instruction data. In fact, this supervised fine-tuning paradigm does not fully exploit the rich information available in the visual modality data or the internal knowledge of frozen large language models (LLMs), yet merely increasing the training data without optimizing data utilization can be inefficient. For example, in conventional video question-answering (VideoQA) tasks, models typically predict answers based on a given video and its associated annotated question. This approach, however, is suboptimal for effective learning. On one hand, as noted in (McQuivey, 2008) that “A Video Is Worth 1.8 Million Words”, videos often contain extensive information that can be described in various forms of language. However, typical datasets offer text that is both limited in length and uniform in structure, which significantly underutilizes the rich information embedded in videos. This restricts the model’s learning to the specific annotated question-answer pairs, limiting its ability to attend to the whole video and generalize to semantically similar questions. Consequently, this naive training paradigm hinders the model’s learning efficiency. On the other hand, this mechanical and passive supervised training method is notably inferior compared to human learning processes, which tends to be more active. Humans often draw upon past experiences and cognition to enrich their understanding of current events, hence proactively pose new questions and seeking answers for a more comprehensive and profound grasp of the situation, bringing higher learning efficiency by knowledge reusing.
@@ 1 Introduction
To address this, we introduce Bootstrapping Video-Language Alignment (BoViLA) training framework via LLM-based self-questioning and answering. It further exploits rich information in the videos by reusing internal knowledge of LLM itself. BoViLA includes two roles, the questioner and the answerer, both played by the same model and improve each other alternately through self-questioning and answering, as shown in Fig. 1. Questioner generates new questions for enabling itself to further extract aligned knowledge from videos and unleash power of the LLMs. Answerer provides feedback to questioner in terms of the self-generated questions. This framework features efficiently employment of video data and the internal historical knowledge of LLMs.
However, especially in the early stages of training, inadequate modal alignment often causes the questioner to generate low-quality questions. This makes training process unstable and may damage the final training results. To address this issue, we introduce an uncertainty estimation method based on evidential deep learning (EDL) to evaluate the uncertainty of LLMs’ generation. Then, we control the influence our self-generated questions based on measured uncertainty. To effectively apply EDL to LLMs, whose most parameters are frozen and logits have high dimensions, we employ a technique called “evidence decoupling” to prevent information loss and ensure the effectiveness of uncertainty estimation.
We verify the effectiveness of the BoViLA on five challenging VideoQA benchmarks: STAR (Wu et al., 2024a), How2QA (Li et al., 2020), DramaQA (Choi et al., 2021), TVQA (Lei et al., 2018), and VLEP (Lei et al., 2020), where BoViLA outperforms several strong baselines. Moreover, we present extensive ablation studies as shown in Table 3. To sum up, our contributions are as follows:
  • 1.
@@
2 Related Work
2.1 LLMs for multi-modal understanding
As LLMs have demonstrated impressive capabilities (Brown et al., 2020; Ouyang et al., 2022; Touvron et al., 2023), there has been increasing interest in exploring multi-modal language models (MLLMs) with visual capabilities. Unlike the more costly joint visual-language pre-training methods, some approaches focus on training lightweight visual-language connectors that endow LLMs with visual abilities (Ma et al., 2023). These methods efficiently utilize the linguistic knowledge of LLMs and the visual knowledge of pre-trained visual encoders for cross-modal alignment. For instance, Flamingo (Alayrac et al., 2022) utilize a cross-attention mechanism to inject visual knowledge into the LLM, which is the so-called Perceiver Resampler. LLaMA-Adapter (Zhang et al., 2023a) applies a linear projection along with prompt adaptation to incorporate visual information, effectively projecting visual embeddings into the input space of the LLM. Additionally, BLIP-2 (Li et al., 2023) trains a module called the Q-former to bridge the modal gap between pre-trained visual encoders and LLMs, enhancing the model’s multi-modal understanding.
2.2 Video Question-Answering
@@
Traditional VideoQA methods relied on training separate visual and text encoders, along with temporal modeling and answering modules (Qian et al., 2023; Xiao et al., 2022; Lei et al., 2021). However, with the advent of large language models (LLMs), there is a growing trend towards using LLM-based approaches due to their advanced reasoning abilities (Yu et al., 2024a; Wang et al., 2023b; Ko et al., 2023; Yu et al., 2024b). For instance, SeViLA (Yu et al., 2024a) uses an LLM to select keyframes from a video and employs another LLM to answer questions based on these keyframes, fully leveraging LLMs’ capabilities. VLAP (Wang et al., 2023b) improves on this by introducing a Frame-Prompter and QFormer-Distiller for more efficient modality alignment. CREMA (Yu et al., 2024b) trains a multimodal Q-former to integrate information from different modalities to answer questions. We notice the recent work of LLaMA-VQA (Ko et al., 2023) as closest to ours, which trains only a linear layer and adaptor to enable LLMs to understand videos, and enhancing modality alignment through multi-task learning by reconstructing both questions and video. In contrast, our work prompts LLMs to simultaneously engage in self-questioning and answering. The major differences are that (i) We use self-generated questions as augmented training data and ask the model to answer them correctly. (ii) We further enhance the model’s questioning ability by encouraging it to answer correctly from self-generated questions. In other words, the loss of the answerer on the self-generated questions propagates gradients back to the parameters of the questioner.
2.3 LLM-Based Bootstrapping Training
@@
Early research has extensively explored the use of LLMs to generate training data for other models, which leverages the capabilities of LLMs to reduce the dependency on human labor for data collection (Lee et al., 2024; Huang et al., 2022; Wang et al., 2024). Recently, there has been increasing interest in using LLMs to generate data for their own training (Ulmer et al., 2024; Zhao et al., 2024; Wang et al., 2022a). For example, (Wang et al., 2022a) enables models to generate data for instruction fine-tuning, focusing on data efficiency and general-purpose tasks. (Huang et al., 2022) use Chain-of-Thought prompting and self-consistency to generate rationale-augmented answers without labeled data. (Ulmer et al., 2024) focuses on a single improve step and employs a conceptually simpler supervised finetuning strategy instead of RL. (Zhao et al., 2024) investigates how self-generation can further enhance an instruction-finetuned model’s ability to execute task-specific instructions. Our work differs from prior efforts in several key ways: (i) We focus on multi-modal tasks rather than text-only ones; (ii) We update model parameters end-to-end rather than alternately performing these processes offline to improve modality alignment; (iii) This dual approach enhances learning efficiency, offering a faster and more convenient solution.
@@
Furthermore, training data generation can be seen as a form of knowledge distillation (Wang et al., 2022a; Yang et al., 2024), while our self-questioning and answering method can also be seen as self-distillation.
2.4 Uncertainty Estimation Model
@@
Numerous studies have explored models capable of estimating uncertainty to enhance reliability and trustworthiness (Amini et al., 2020; Sensoy et al., 2018). EDL (Sensoy et al., 2018), as one of these approaches, models ”second-order probabilities” over logits based on Dempster-Shafer Theory (Shafer, 1992) and Subjective Logic (Jsang, 2018) to capture uncertainty conveniently and accurately in various fields (Shao et al., 2024; Holmquist et al., 2023; Huang et al., 2024), particularly for out-of-distribution (OOD) samples. In this work, we leverage EDL-estimated uncertainty to evaluate the modality alignment within the context and employ it for ”soft filtering” of the model’s self-generated questions. To the best of our knowledge, we are the first to explore the integration of EDL with LLMs.
3 Methdology
@@ 3.1 Model Architecture
As shown in Fig. 2, our model architecture consists of an LLM decoder, a learnable linear layer for mapping visual tokens to the text embedding space, a lightweight adaptor for task-specific fine-tuning, and an EDL head to estimate uncertainty. For videos, we first extract frames $\bm{v}=\{v_{1},v_{2},\cdots,v_{N_{v}}\}$ and use a pretrained visual encoder $E(\cdot)$ to extract their features $E(\bm{v})=\{E(v_{1}),E(v_{2}),\cdots,E(v_{N_{v}})\}\in{\mathbb{R}}^{{N_{v}}\times D}$. These features are then mapped to the text embedding space with the learnable linear layer $f_{\theta}(\cdot)$, and an extra learnable temporal embedding $\bm{t}=\{t_{1},t_{2},\cdots,t_{N_{v}}\}\in{\mathbb{R}}^{{N_{v}}\times D}$ is added, i.e.
@@
For text inputs such as task instructions, questions, or answers, we use the LLM’s tokenizer and token embedding module to obtain the corresponding tokens and embedding. Specifically, for questions, we get the tokens $\bm{q}=\{q_{1},q_{2},\cdots,q_{N_{q}}\}$ and their embedding $\bm{h}^{0}_{q}=\{q^{0}_{1},q^{0}_{2},\cdots,q^{0}_{N_{q}}\}$. Similarly, for answers, we obtain the tokens $\bm{a}=\{a_{1},a_{2},\cdots,a_{N_{a}}\}$ and their embedding $\bm{h}^{0}_{a}=\{a^{0}_{1},a^{0}_{2},\cdots,a^{0}_{N_{a}}\}$. Then, we concatenate the video and text tokens together as input to the LLM. As task-specific fine-tuning is necessary, we employed several prevalent PEFT methods such as LoRA (Hu et al., 2021), adapter methods (Hu et al., 2023; Zhang et al., 2023b), and prefix-tuning (Li and Liang, 2021; Zhang et al., 2023b) for efficient modal alignment.
3.2 Bootstrapping Training Framework
@@
3.2.1 Questioner.
When the model acts as the questioner, we instruct it to generate a question based on the video, answer, and seed question, as shown in Figure 2. Assuming the LLM has $l$ layers, the hidden states from the final layer are $h^{l}_{q}=\{q^{l}_{1},q^{l}_{2},\cdots,q^{l}_{N_{q}}\}$. For gradient backpropagation, we use Gumbel-Softmax (Jang et al., 2016) for sampling, formulated as below:
@@
$\mathcal{L}_{\mathrm{v\overline{q}a}}=-\log p(\bm{a}|\bm{v},\bm{\overline{q}})=-\sum\nolimits_{i=1}^{N_{a}}\log p(a_{i}|\bm{v},\bm{\overline{q}},a_{<i}).$ (6)
We provide the detailed prompt template of BoViLA in Table 1.
Table 1: Input prompt of questioner and answerer.
3.2.3 Regularization Based on Seed Questions.
A key point of BoViLA is the differentiability of self-generated questions. If answerer fails to predict the target answer from $\overline{q}$, then penalty will be applied to the questioner to improve itself. In contrast, if answerer succeed in answering $\overline{q}$, that means $\overline{q}$ is considered a ”good” question by answerer. However, this can easily lead to ”information leakage”, where the questioner creates meaningless questions but contain target answers implicitly. To address this, we apply seed-based question regularization to constrain the generation space of questioner as follows:
@@
Detailed ablation experiments are presented in Section 3, qualitatively and quantitatively demonstrating the effectiveness of each design, including the trainging with augmented question-answer samples and the regularization trick.
3.3 EDL-based Filter